# FLM-Audio: Natural Monologues Improves Native Full-Duplex Chatbots via Dual Training

### v2

@@ 3.2 Stage-2: Supervised Fine-tuning (SFT)
Supervised Fine-tuning (SFT) is applied to incorporate the capabilities to function as a general-purpose SDM. In FLM-Audio, we set up two sub-stages, including a semi-duplex Fine-tuning-1, followed by a final stage Fine-tuning-2 (Table 2).
@@ 3.2.1 Data Collection
Transcript Collection. We curate textual Chinese and English instruct-following data from open-source corpora, including Magpie magpie , Belle belle , Infinity-Instruct infinity ; infinityinstruct , WizardLM wizardlm , and Ultrachat ultrachat . User instructions are retained, while responses are refined using the DeepSeek-V3 deepseek-v3 API. To ensure suitability for TTS, we enforce constraints on length, style, and the use of special symbols. Dialog lengths vary from 1 to 10 turns, mixing natural multi-turn conversations (e.g., Ultrachat) with synthesized single-turn instruct-following examples. In total, we sample 200K dialogs as transcripts for speech synthesis.
TTS Generation. We collect over 700 human voices, filter them based on DNSMOS dnsmos , and use the selected voices as references for a locally deployed Fishaudio TTS system fishspeech . For each textual transcript, two distinct user voices are sampled, while the model’s voice remains consistent across all dialogs.
Noise Augmentation. To improve robustness, we augment training audio with both environmental and speech noise. Sources include the DNS Challenge dataset dnschallenge , RNNoise [^8] https://github.com/xiph/rnnoise , and random speech clips from Stage-1 post-training data. For each training sample, we add concatenated random noise segments to the listening channel waveforms. With probability 0.6, wave gain is applied to user utterances, scaling amplitudes within a range of -24 to +20 dB. We enforce a minimum final loudness of -40 dB.  [^9] We compute $dB=20.0\times\log_{10}(\text{wave\_root\_mean\_square}+1e{-6})$. Noise clips are randomly scaled to (-70, -40) dB. Additionally, with probability 0.3, noise segments are replaced with silence.
Table 3: Training configuration for different stages. All the learning rate decay follow a cosine schedule.
@@ 3.2.2 Sub-stages
FLM-Audio adopts a fundamentally different SFT sub-stage arrangement from Moshi moshi . We first introduce a semi-duplex transition stage, Fine-tuning-1, which is designed to combine the TTS- and ASR-alignment capabilities learned during post-training. The token streams are organized as follows:
ASR-Response-TTS Style. As illustrated in Figure 3 (middle), utterances are arranged in a semi-duplex manner. The model first processes the entire user instruction and immediately transcribes it into ASR tokens in the monologue channel. This span begins with a special <asr> token and terminates with an <answer> token. During the ASR phase, the speaking channel remains silent. Once the <answer> token is reached, the model generates a textual response, and with a delay of 2 steps, produces the corresponding speech output (a TTS rendering of the response) on the speaking channel. Following Moshi, a 1-step offset is maintained between the semantic channel and the seven acoustic channels.
This format effectively combines the TTS-style and ASR-style training signals from Stage-1, embedding both capabilities into each dialog instance and facilitating smooth transfer between post-training and SFT.
After this transitional stage, we proceed to the final Fine-tuning-2, which uses the same dialog transcripts but reorganized to hide the ASR supervision:
Response-TTS Style. As shown in Figure 3 (top right), we remove the ASR text from the semi-duplex ASR-Response-TTS format, retaining only the response monologue. In this setting, the model is required to infer the user’s intent directly from audio input and generate the appropriate textual and spoken responses. After this stage, FLM-Audio achieves response latency equivalent to Moshi, while maintaining strong language modeling performance.
Response-TTS Style with Free User Inputs. As shown in Figure 3 (bottom right), we further break the semi-duplex structure to enable full duplexity. Here, user utterances may occur at arbitrary times, potentially interrupting the model’s response, forcing the model to learn realistic turn-taking. Specifically, when interrupted by meaningful user speech, the model must cut off both its monologue and speaking channels, falling silent within a short delay. Once the interruption ends, it resumes dialog generation, potentially addressing a new topic. To simulate this behavior, interruptions are introduced with probability 0.7, and the reaction delay is tuned to 0.5 seconds to avoid oversensitivity to short back-channels.
Following Moshi, we also apply speech leakage augmentation by mixing the speaking channel back into the listening channel with probability 0.3, applying a random gain (0-0.2) and a random delay (0.1-0.5 seconds) to enhance robustness in microphone-based interaction. After training on this data, FLM-Audio becomes a fully duplex, responsive audio chatbot.
@@ 4 Experiments
As discussed above, FLM-Audio features native full-duplex design with natural monologues, and a 4-stage training paradigm with dual formats for data organization. Thus, we focus on answering the following three research questions with experimental observations:
  • RQ1: In native full-duplex systems, do natural monologues improve semantic understanding as hypothesized?
  • RQ2: How effective is the dual data-format strategy across training stages, and how crucial is it to final performance?
  • RQ3: How does FLM-Audio compare against state-of-the-art full-duplex chatbots in terms of responsiveness, speech quality, and dialog capability?
To address these questions, we benchmark FLM-Audio against representative clusters of existing models and systems across three dimensions: audio understanding, audio generation, and duplex dialog performance. In addition, we conduct ablation studies under different training configurations to isolate the effects of natural monologues and dual-format supervision.
@@ 4.1 Audio Understanding
We evaluate audio understanding through automatic speech recognition (ASR) and spoken question answering tasks. For ASR, we adopt word error rate (WER) as the primary metric, testing on both Chinese and English benchmarks, including Fleurs-zh fleurs and LibriSpeech-clean librispeech . While instruction-following with spoken input is addressed separately in Section 4.3, we also include LlamaQuestions llamaquestions as a speech-based QA benchmark, reporting accuracy.
For comparison, we include Whisper-large-v3 whisper , Qwen2-Audio qwen-audio , MinMo minmo , and GLM-4-Voice glm-voice , all of which are specialized audio–language models, as well as GPT-4o 4o , a proprietary large-scale system.
@@
zh clean
GPT-4o 5.4 - 71.7
Whisper-large-v3 7.7 1.8 -
Qwen2-Audio 7.5 1.6 -
MinMo 3.0 1.7 64.1
GLM-4-Voice - 2.8 50.0 (64.7)
@@
We emphasize the comparison to Moshi moshi , the only other native full-duplex audio–language model. Despite being trained with less than 15% of Moshi’s audio data and without fine-grained timestamps, FLM-Audio achieves superior performance: on LibriSpeech-clean, FLM-Audio yields significantly lower WER. Furthermore, whereas Moshi is specialized for English, more than half of FLM-Audio ’s training data is Chinese, enabling broader multilingual coverage.
Finally, we note a pronounced improvement in Chinese ASR performance after the Post-Training-2 stage. This aligns with our training setup, where Post-Training-2 replaces coarse ASR annotations with high-quality, human-annotated Chinese transcripts. English ASR, by contrast, already performs competitively after Post-Training-1 even without additional fine annotations, suggesting that our natural monologue design provides a key advantage for capturing audio semantics.
@@ 4.2 Audio Generation
We assess audio generation performance using the Seed-TTS-en and Seed-TTS-zh benchmarks seed-eval , following the standard evaluation protocols. Results are presented in Table 5.
While FLM-Audio is not explicitly optimized for high-fidelity voice cloning-and therefore does not surpass state-of-the-art TTS systems in similarity (SIM) scores-it achieves word error rate (WER) performance comparable to advanced, specialized TTS models such as Seed-TTS seed-tts and CosyVoice cosyvoice . Moreover, its WER scores are also on par with those of general audio–language models, including GLM-4-Voice and MinMo.
@@ 4.3 Full-duplex Chatting
For LLM-based assistants, full-duplex chatting differs substantially from traditional text-based multimodal instruction-following, particularly with respect to human preference. In instruction-following tasks, users often value detailed, elaborate responses, such as those required for programming or complex reasoning o1 ; dpskr1 . In natural spoken conversations, however, users typically prefer concise, summarized, or even intentionally evasive replies.
To capture these differences, we conduct a comprehensive evaluation combining both automatic metrics and human judgment.
Automatic evaluation.
We construct a speech instruction-following test set using publicly available Chinese prompts formatted in the style of AlpacaEval alpaca_eval . Prompts are converted into audio using our TTS pipeline. DeepSeek-V3 deepseek-v3 is employed as a reference model to assign quality scores (0–10 scale) by comparing candidate textual responses to ground-truth answers.
Human evaluation.
We run a double-blind, head-to-head comparison between FLM-Audio and Qwen2.5-Omni qwen-omni , a state-of-the-art streaming chatbot. Five human annotators rate multi-turn audio responses across four dimensions: (1) Helpfulness, standing for the informativeness and relevance of content; (2) Naturalness, for conversational tone and linguistic fluency; (3) Responsiveness, representing reaction speed to interruptions and dynamic user input; and (4) Robustness, which means stability under noisy real-world conditions. This benchmark shares the same spirit as sdm-bench , but is constructed in Chinese.
Results.
Table 6 summarizes the results. Compared with Qwen2.5-Omni, FLM-Audio delivers responses of comparable quality in terms of helpfulness, as confirmed by both automatic scoring and human ratings. More importantly, in dimensions that matter most for real-time interaction: naturalness, responsiveness, and robustness, FLM-Audio demonstrates a clear advantage. We attribute this to the model’s native full-duplex design and the effectiveness of the dual training paradigm.
Ablation study.
We further compare against a baseline trained without the semi-duplex Fine-tuning-1 stage (i.e., omitting ASR-style supervision). This variant shows a marked drop in instruction-following ability, underscoring the importance of retaining the dual data format in SFT. In particular, the ASR-style organization significantly strengthens audio understanding, validating the design of our training pipeline.
@@ Results.
4.4 Answers to Research Questions
We now revisit the research questions posed at the beginning of Section 4:
RQ1: In native full-duplex systems, do natural monologues improve semantic understanding as hypothesized? Yes. As shown in Section 4.1, FLM-Audio matches Moshi’s performance after only the Post-training-1 stage, despite being trained with less than 15% of Moshi’s audio data. Since both models share the same RQ-Transformer backbone and rely on coarse third-party ASR transcripts at this stage, the performance advantage is best explained by our natural monologue design. Additional evidence comes from final instruct-following results: FLM-Audio reaches performance levels comparable to state-of-the-art systems such as Qwen-2.5-Omni, whereas Moshi has been reported to lag behind in this area omniflatten . These results also support our hypothesis that meaningful modifications in data organization, alignment strategies, and training paradigms can bring significant gains with the same LLM model architecture kept intact.
RQ2: How effective is the dual data-format strategy across training stages, and how crucial is it to final performance? The TTS-style format is essential for any responsive full-duplex system, including both FLM-Audio and Moshi. Thus, the key question is whether the additional ASR-style format provides a measurable benefit. Results in Table 4 and Table 6 confirm that it does: models trained without ASR-style supervision show clear disadvantages in audio understanding and instruction-following, underscoring the importance of dual-format training.
RQ3: How does FLM-Audio compare against state-of-the-art full-duplex chatbots in terms of responsiveness, speech quality, and dialog capability? As demonstrated in Section 4.3, FLM-Audio achieves comparable overall response quality to Qwen-2.5-Omni, while delivering superior naturalness, responsiveness, and robustness in interactive settings. These results affirm that native full-duplex design, coupled with our dual training paradigm, enhances both the quality and the real-time usability of spoken dialog systems.
5 Conclusion and Future Challenges
In this paper, we explored ways to advance native full-duplex audio chatbots. We introduced a novel alignment strategy, natural monologues, together with a training pipeline that integrates dual ASR-like and TTS-like capabilities. Building on this design, we developed and released FLM-Audio, a bilingual chatbot with native full duplexity. Compared with the most related baseline Moshi, FLM-Audio achieves equivalent response latency while delivering substantially stronger language modeling and instruction-following performance. Our experimental results validate the effectiveness of natural monologues and support our central hypothesis that they benefit native full-duplex systems when combined with dual training.
@@ 5 Conclusion and Future Challenges
Constrained by data volume and computational resources, we have not yet scaled FLM-Audio to larger parameter counts—a direction where native duplex models could exhibit even greater advantages over TDM-based approaches. We hope this research will inspire further exploration into scaling native full-duplex architectures, both to push the performance upper bound of task-solving and to provide a more comprehensive comparison against TDM-based solutions.

### v3

@@ 4.1 Stage-2: Supervised Fine-tuning (SFT)
Supervised Fine-tuning (SFT) is applied to incorporate the capabilities for a general-purpose SDM. In FLM-Audio, we set up two sub-stages, including a semi-duplex Fine-tuning-1, followed by a final stage Fine-tuning-2 (Tables 2, 8).
@@ 4.1.1 Data Collection
Transcript Collection. We curate textual Chinese and English instruct-following data from open-source corpora, including Magpie (Xu et al., 2024b), Belle (Ji et al., 2023), Infinity-Instruct (Zhao et al., 2024; Li et al., 2025), WizardLM (Xu et al., 2023), and Ultrachat (Ding et al., 2023). User instructions are retained, while responses are refined using the DeepSeek-V3 (Liu et al., 2024) API. To ensure suitability for TTS, we enforce constraints on length, style, and the use of special symbols. Dialog lengths vary from 1 to 10 turns, mixing natural multi-turn conversations (e.g., Ultrachat) with synthesized single-turn instruct-following examples. In total, we sample 200K dialogs as transcripts for speech synthesis.
Audio Generation. We collect over 700 human voices, filter them based on DNSMOS (Reddy et al., 2022), and use the selected voices as references for a locally deployed Fishaudio TTS system (Liao et al., 2024). For each textual transcript, two distinct user voices are sampled, while the model’s voice remains consistent across all dialogs. To improve robustness, we augment training audio with both environmental and speech noise. The processing details are provided in Appendix B.
Table 2: Training configuration for different stages. The learning rate follows cosine schedules.
@@ 4.1.2 Sub-stages
For SFT, we first introduce a semi-duplex transition stage, Fine-tuning-1, which integrates the TTS and ASR capabilities learned during post-training. The token streams are organized as follows:
ASR-Response-TTS Style. As illustrated in Figure 3 (middle), utterances are arranged in a semi-duplex manner. The model first processes the entire user instruction and immediately transcribes it into ASR tokens in the monologue channel. This span begins with a special <asr> token and terminates with an <answer> token. During the ASR phase, the speaking channel remains silent. Once the <answer> token is reached, the model generates a textual response, and, with a delay of 2 steps, produces the corresponding speech output (a TTS rendering of the response) on the speaking channel. Following Moshi, a 1-step offset is maintained between the semantic channel and the seven acoustic channels. This style of data effectively combines the TTS-style and ASR-style training signals from Stage-1, embedding both capabilities into each dialog instance and facilitating smooth transfer between post-training and SFT.
After this transitional stage, we proceed to the final stage, Fine-tuning-2, which uses the same dialog transcripts but reorganizes the textual and/or audio tokens:
Response-TTS Style. As shown in Figure 3 (top right), we remove the ASR text from the semi-duplex ASR-Response-TTS format, retaining only the response monologue. In this setting, the model is required to infer the user’s intent directly from audio input and generate the appropriate textual and spoken responses. After this stage, FLM-Audio achieves a response latency equivalent to Moshi, while maintaining strong language modeling performance.
Response-TTS Style with Free User Inputs. As shown in Figure 3 (bottom right), this style enables full duplexity. Here, user utterances may occur at arbitrary time, potentially interrupting the model’s response, forcing the model to learn realistic turn-taking. Specifically, when interrupted by meaningful user speech, the model must cut off both its monologue and speaking channels, falling silent within a short delay. Once the interruption ends, it resumes dialog generation, potentially addressing a new topic. To simulate this behavior, interruptions are introduced with probability 0.7, and the reaction delay is tuned to 0.5 seconds to avoid oversensitivity to short back-channels.
@@ 5 Experiments
As discussed above, FLM-Audio features native full-duplex design with contiguous monologues, and a 4-stage training paradigm with dual formats for data organization. Thus, we focus on answering the following three research questions with experimental observations: RQ1: In native full-duplex systems, do contiguous monologues improve semantic understanding as hypothesized? RQ2: How effective is the dual data-format strategy across training stages, and how crucial is it to final performance? RQ3: How does FLM-Audio compare against state-of-the-art full-duplex chatbots in terms of responsiveness, speech quality, and dialog capability? To address these questions, we benchmark FLM-Audio against representative clusters of existing models and systems across three dimensions: audio understanding, audio generation, and duplex dialog performance. In addition, we conduct ablation studies under different training configurations to isolate the effects of contiguous monologues and dual-format supervision.
@@ 5.1 Audio Understanding
We evaluate audio understanding through automatic speech recognition (ASR) and spoken question answering tasks. For ASR, we adopt word error rate (WER) as the primary metric, testing on both Chinese and English benchmarks, including Fleurs-zh (Conneau et al., 2022) and LibriSpeech-clean (Panayotov et al., 2015). While instruction-following with spoken input is addressed separately in Section 5.3, we also include LlamaQuestions (Nachmani et al., 2023) as a speech-based QA benchmark, reporting accuracy.
For comparison, we include Whisper-large-v3 (Radford et al., 2023), Qwen2-Audio (Chu et al., 2023), MinMo (Chen et al., 2025), and GLM-4-Voice (Zeng et al., 2024), all of which are specialized audio–language models, as well as GPT-4o (Hurst et al., 2024), a proprietary large-scale system.
@@
zh (WER $\downarrow$) clean (WER $\downarrow$) (Acc. $\uparrow$)
GPT-4o (Hurst et al., 2024) 5.4 - 71.7
Whisper-large-v3 (Radford et al., 2023) 7.7 1.8 -
Qwen2-Audio (Chu et al., 2023) 7.5 1.6 -
MinMo (Chen et al., 2025) 3.0 1.7 64.1
Freeze-Omni (Wang et al., 2024b) - 3.82 72
OmniFlatten (Zhang et al., 2024a) - 7.91 -
GLM-4-Voice (Zeng et al., 2024) - 2.8 50.0 (64.7)
Kimi-Audio (Ding et al., 2025) 2.69 1.28 -
Qwen-2.5-Omni (Xu et al., 2025) 2.92 2.37 -
@@
We emphasize the comparison to Moshi (Défossez et al., 2024), the only other native full-duplex audio–language model. Despite being trained with less than 15% of Moshi’s audio data and without fine-grained timestamps, FLM-Audio achieves superior performance: on LibriSpeech-clean, FLM-Audio yields significantly lower WER. Furthermore, whereas Moshi is specialized for English, more than half of FLM-Audio ’s training data is Chinese, enabling broader multilingual coverage.
Finally, we note a pronounced improvement in Chinese ASR performance after the Post-Training-2 stage. This aligns with our training setup, where Post-Training-2 replaces coarse ASR annotations with high-quality, human-annotated Chinese transcripts. English ASR, by contrast, already performs competitively after Post-Training-1 even without additional fine annotations, suggesting that our contiguous monologue design provides a key advantage for capturing audio semantics.
@@ 5.2 Audio Generation
We assess audio generation performance using the Seed-TTS-en and Seed-TTS-zh benchmarks (Anastassiou et al., 2024a), following the standard evaluation protocols. Results are presented in Table 4.
While FLM-Audio is not explicitly optimized for high-fidelity voice cloning-and therefore does not surpass state-of-the-art TTS systems in similarity (SIM) scores-it achieves word error rate (WER) performance comparable to advanced, specialized TTS models such as Seed-TTS (Anastassiou et al., 2024b) and CosyVoice (Du et al., 2024). Moreover, its WER scores are also on par with those of general audio–language models, including GLM-4-Voice and MinMo.
@@ 5.3 Full-duplex Chatting
For LLM-based assistants, full-duplex chatting differs substantially from traditional text-based multimodal instruction-following, particularly with respect to human preference. In instruction-following tasks, users often value detailed, elaborate responses, such as those required for programming or complex reasoning (Jaech et al., 2024; DeepSeek-AI et al., 2025). In natural spoken conversations, however, users typically prefer concise, summarized, or even intentionally evasive replies. To capture these differences, we conduct a comprehensive evaluation combining both automatic metrics and human judgment.
@@
Textual Automatic evaluation.
We construct a speech instruction-following test set using publicly available Chinese prompts formatted in the style of AlpacaEval (Li et al., 2023). Prompts are converted into audio using our TTS pipeline. DeepSeek-V3 (Liu et al., 2024) is employed as a reference model to assign quality scores (0–10 scale) by comparing candidate textual responses to ground-truth answers.
Table 6: Objective evaluation on noisy test set with interruptions.
Turn-taking Acc. $\uparrow$ $|\Delta\text{Pause}|$ $\downarrow$ $|\Delta\text{Gap}|$ $\downarrow$ $|\Delta\text{Overlap}|$ $\downarrow$
0.98 1.9 0.9 2.3
Objective Evaluation for Full-duplexity
We additionally compute objective metrics on a test set containing background noise and random user interruptions. The reported metrics include turn-taking accuracy and the per-minute deviation from ground-truth in pause, gap, and overlap durations, following the definitions in (Wang et al., 2025; Nguyen et al., 2023). As in most prior work, the training and test data are drawn from the same distribution, and these objective statistics quantify how effectively the model learns full-duplex dialog behaviors. Results are presented in Table 6.
Human Evaluation.
We run a double-blind comparison between FLM-Audio and Qwen2.5-Omni (Xu et al., 2025), a state-of-the-art streaming chatbot. Five human annotators rate multi-turn audio responses across four dimensions: (1) Helpfulness, standing for the informativeness and relevance of content; (2) Naturalness, for conversational tone and linguistic fluency; (3) Responsiveness, representing reaction speed to interruptions and dynamic user input; and (4) Robustness, which means stability under noisy real-world conditions. This benchmark shares the same spirit as (Lin et al., 2025), but is constructed in Chinese.
Results.
Table 5 summarizes the results. Compared with Qwen2.5-Omni, FLM-Audio delivers responses of comparable quality in terms of helpfulness, as confirmed by both automatic scoring and human ratings. More importantly, in dimensions that matter most for real-time interaction: naturalness, responsiveness, and robustness, FLM-Audio demonstrates a clear advantage. We attribute this to the model’s native full-duplex design and the effectiveness of the dual training paradigm. We further compare against a baseline trained without the semi-duplex Fine-tuning-1 stage (i.e., omitting ASR-style supervision). This variant shows a marked drop in instruction-following ability, underscoring the importance of retaining the dual data format in SFT. In particular, the ASR-style organization significantly strengthens audio understanding, validating the design of our training pipeline.
5.4 Controlled Comparison: Contiguous vs. Word-level
Because FLM-Audio relies on contiguous monologues beginning from the first post-training stage, conducting a fully controlled comparison against a word-level alignment strategy trained on the entire dataset would be prohibitively expensive in terms of computational resources. To provide a quick but informative sanity check, we perform an ablation-style comparison by processing 5% of our post-training data using a Moshi-like (Défossez et al., 2024) word-level alignment method (ASR-style, where each word-level token slightly lags behind its pronunciation). We train this Moshi-like model under the same configuration and run a complete pass over this 5% subset. At the end of training, we evaluate both models using ASR word error rate (WER) on LibriSpeech-clean and acc_norm on HellaSwag (Zellers et al., 2019). Results are shown in Table 7. We find that our contiguous strategy converges substantially faster than the word-level alternative, consistent with our observation that FLM-Audio achieves comparable or better performance with significantly less training data than Moshi.
Table 7: Ablation analysis: contiguous monologues vs. word-level alignment.
Strategy Fleurs-zh (WER $\downarrow$) HellaSwag (Acc. $\uparrow$)
Contiguous (FLM-Audio) 18.2 61.6
Word-Level (Moshi-like) 22.3 58.3
6 Conclusion
@@
In this paper, we introduced contiguous monologues for native full-duplex audio-language models, together with a dual training pipeline that integrates ASR- and TTS-like capabilities. Building on this design, we developed FLM-Audio, a bilingual chatbot. Compared with the most related baseline Moshi, FLM-Audio achieves equivalent response latency while delivering substantially stronger language modeling performance with less data. It also outperforms TDM-based systems in dialog experiences. Constrained by data volume and computational resources, we have not yet scaled FLM-Audio to larger parameter counts—a direction where native duplex models could exhibit even greater advantages over TDM-based approaches. We hope this research will inspire further exploration into scaling native full-duplex architectures, both to push the performance upper bound of task-solving and to provide a more comprehensive comparison against TDM-based solutions.
ETHICS STATEMENT
The data used to train FLM-Audio is obtained exclusively from publicly available sources or through commercial licenses. No unauthorized or private data has been included. As FLM-Audio is developed upon a foundation language model and refined through post-training, harmful contents could potentially be elicited from the released model despite the efforts made for safety. The generated contents by FLM-Audio do not represent the opinions of the authors or entities involved.
@@ Acknowledgments
Appendix A Training stage comparison with Moshi
We summarize different training stages compared to Moshi in Table 8. Both models undergo four stages in total, including two post-training and two fine-tuning stages. However, to better exploit the language modeling benefits of contiguous monologues, FLM-Audio features special designs to enhance sentence-level alignment with both listen and speak channels during the early stages.
Table 8: Training Paradigm: FLM-Audio and Moshi.
Moshi (Défossez et al., 2024) Helium 1-channel 2-channel semi-duplex full-duplex dialog full-duplex instruct
FLM-Audio Qwen-2.5-vl 2-channel coarse 2-channel fine semi-duplex w/ ASR full-duplex w/o ASR
Appendix B Noise Augmentation for SFT
Noise sources include the DNS Challenge dataset (Dubey et al., 2023), RNNoise [^7] https://github.com/xiph/rnnoise , and random speech clips from Stage-1 post-training data. For each training sample, we add concatenated random noise segments to the listening channel waveforms. With probability 0.6, wave gain is applied to user utterances, scaling amplitudes within a range of -24 to +20 dB. We enforce a minimum final loudness of -40 dB. We compute $dB=20.0\times\log_{10}(\text{wave\_root\_mean\_square}+1e{-6})$. Noise clips are randomly scaled to (-70, -40) dB. Additionally, with probability 0.3, noise segments are replaced with silence.
Following Moshi, in the final stage, we also apply speech leakage augmentation by mixing the speaking channel back into the listening channel with probability 0.3, applying a random gain (0-0.2) and a random delay (0.1-0.5 seconds) to enhance robustness in microphone-based interaction.